HETAV PATEL

Grade 11,

the sides of a triangle are 2x-y+5=0,x+y-5=0,x-2y-5=0 sum of tangents of its interior angles is ______________________

the sides of a triangle are 2x-y+5=0,x+y-5=0,x-2y-5=0 sum of tangents of its interior angles is ______________________

Grade:12th Pass

2 Answers

Rishabh Raj
12 years ago

At one vertex, 2x-y+5=0 and x+y-5=0 meet so by using simultaneous equations x=0 and y=5. Thus the co-ordinates of that vertex is [0,5]. By using this method the other vertex wud be [-5,-5] and [5.,0]. Now find the length of each side by distance formula and the length will be sqrt125,sqrt125 and sqrt50. Now, draw a perpendicular line from a vertex to the side of length sqrt50. The perp. line wud bisect that line as the. Now u can find the tangents of the angles.

Aman Bansal
12 years ago

Dear Archu,

At one vertex, 2x-y+5=0 and x+y-5=0 meet so by using simultaneous equations x=0 and y=5. Thus the co-ordinates of that vertex is [0,5]. By using this method the other vertex wud be [-5,-5] and [5.,0]. Now find the length of each side by distance formula and the length will be sqrt125,sqrt125 and sqrt50. Now, draw a perpendicular line from a vertex to the side of length sqrt50. The perp. line wud bisect that line as the. Now u can find the tangents of the angles.

Best Of Luck...!!!!

Cracking IIT just got more exciting,It’s not just all about getting assistance from IITians, alongside Target Achievement and Rewards play an important role. ASKIITIANS has it all for you, wherein you get assistance only from IITians for your preparation and win by answering queries in the discussion forums. Reward points 5 + 15 for all those who upload their pic and download the ASKIITIANS Toolbar, just a simple  to download the toolbar….

So start the brain storming…. become a leader with Elite Expert League ASKIITIANS

Thanks

Aman Bansal

Askiitian Expert